# Constrained Stein Variational Trajectory Optimization

### v2

@@ VII-A7 Initialization
As introduced in section IV, we have a user-specified prior on controls $p(U)$. To initialize CSVTO on a new problem, we proceed by sampling from this prior $p(U)$ and using the dynamics $f(x_{t},u_{t})$ to generate sampled trajectories. In this way we ensure that the initial trajectory satisfies the dynamics constraints.
We use a different initialization scheme when running trajectory optimization online in an MPC style-loop as in Algorithm 2, as it is typical to warm-start the optimization with the solution from the previous timestep. For a single particle, the trajectory consists of $\tau=(x_{1},...,x_{T},u_{0},...,u_{T-1)})$. The shift operation computes $\tau^{\prime}=(x_{2},...,x^{\prime}_{T+1},u_{1},...,u^{\prime}_{T})$. Here $x^{\prime}_{T+1},u^{\prime}_{T}$ is the initialization for the newly considered future timestep. The initializations $x^{\prime}_{T+1},u^{\prime}_{T}$ may be chosen in a problem-specific way. In our approach, we choose them by duplicating the previous timestep’s state and control, i.e. $(x^{\prime}_{T+1},u^{\prime}_{T})=(x_{T},u_{T-1})$.
When running the algorithm with inequality constraints, for both the online and warm-start optimizations we initialize the slack variable $z$ with $z=\sqrt{2|g(\tau)|}$ so that trajectories satisfying the inequality constraint are initialized to satisfy the transformed equality constraint.
@@ VII-A8 Re-sampling
$w_{i}=\frac{\exp(-\frac{\hat{C}_{\lambda}(\hat{tau}_{i})}{\beta})}{\sum^{N}_{j}\exp(-\frac{\hat{C}_{\lambda}(\hat{tau}_{j})}{\beta})},$ (50)
@@
1:function CSVTO($x_{0},\tau,K,\texttt{anneal}$)
2:     $z\leftarrow\sqrt{2|g(\tau)|}$
3:     $\hat{\tau}\leftarrow[\tau,z]^{T}$
@@
14:         $*u_{t}\leftarrow\tau*$
15:         $x_{t}\leftarrow$ StepEnv$(u_{t})$
@@ VIII-A Baselines
We evaluate our method on a 12DoF underactuated quadrotor problem. The goal is to navigate the quadrotor from a start state to a goal state. We chose this problem to demonstrate our approach on a problem with complex nonlinear underactuated dynamics. The experimental setup is shown in Figures 3 and 4. The state of the quadrotor is 6DoF pose parameterized with Euler angles, and the linear and angular velocities. The dynamics for the 12DoF quadrotor are from [64] and are given by
@@
where $c(p),s(p),t(p)$ are $\cos,\sin,\tan$ functions, respectively. We use parameters $m=1kg,I_{x}=0.5kg\cdot m^{2},I_{y}=0.1kg\cdot m^{2},I_{z}=0.3kg\cdot m^{2},K=5,g=-9.81m\cdot s^{-2}$. We use the same dynamics both for planning and for simulation. We place bounds constraints on the $(x,y)$ location of the quadrotor to be within a $10m\times 10m$ area centred at $(0,0)$. The goal is to travel from start locations sampled uniformly from $x,y\in[3.0m,4.5m]$ to a goal location of $(4,4)$ within 100 time steps. We place an equality constraint that the quadrotor must travel along a nonlinear surface $z=f_{surf}(x,y)$. For this surface, we sample $z$ values from a Gaussian Process (GP) prior with an RBF kernel and zero mean function on a $10\times 10$ grid of $(x,y)$ points. We use the sampled values as observations for a GP with the same kernel and mean function, and use the corresponding posterior mean function as the equality constraint. We sample a single surface in this way and use it for all experiments.
In addition, we consider three variants of this task with different obstacle avoidance constraints: 1) We consider the case with no obstacles; 2) We consider the case of static obstacles. For the static obstacles case, we wish to demonstrate our method in a cluttered environment with arbitrarily shaped obstacles. We do this by generating the obstacles similarly to the surface constraint, which results in a smooth obstacle constraint. We consider a constraint function $f_{obs}(x,y)$, where the obstacle-free region is $\{(x,y),f_{obs}(x,y)\leq 0\}$. We sample values for $f_{obs}(x,y)$ from a GP prior with an RBF kernel and a constant mean function of $-0.5$, so that there is a bias towards being obstacle-free, on a $10\times 10$ grid of $(x,y)$ points. We then use these points as the observations for a GP with the same mean function and kernel as the GP prior. We also add observations at $(-4,-4)$ and $(4,4)$ of $-2$, to ensure the start and goal regions are obstacle free. We use the resulting GP posterior mean function as $f_{obs}(x,y)$. We do this once and keep the same obstacle constraint for all trials. The resulting obstacle constraint is shown in Figure 4; 3) Finally, we consider a cylindrical obstacle in the x-y plane that moves during the trial in a path that is unknown to the planner; at every timestep, the planner plans assuming the obstacle will remain fixed. If the quadrotor collides with an obstacle during execution then we consider the task failed.
@@
The control cost is equivalent to the prior on controls $p(u_{t})=\mathcal{N}(0,R^{-1})$. The values we use for the costs are
$Q$ $=\text{Diag}(5,5,0.5,2.5,2.5,0.025,2.5,2.5,2.5,2.5,2.5,2.5)$ (54)
@@
We ran the quadrotor experiments for the three different obstacle cases for 20 trials with randomly sampled starts. The results are shown in Figure 6. CSVTO succeeds for 20/20 trials for the no obstacles and dynamic obstacles cases, and 19/20 for the static obstacle case, all with a goal threshold of 0.3m. For the static-obstacle and dynamic-obstacle experiments, IPOPT is the next best performing with 20/20 trials for no-obstacles at a goal threshold of 0.4m, but success falls to 15/20 for both the static-obstacles and dynamic-obstacle case. We see that running IPOPT with more iterations improves performance for the static obstacles case, but in the other two cases, there is no significant difference in performance when allowing IPOPT to run until convergence. However, running IPOPT to convergence has substantially higher computational time, which we will discuss further in section VIII-E. For the no-obstacles and dynamic-obstacle cases, we see that sample-based methods perform well according to the task success rate, however, they fail to satisfy the surface equality constraint. In addition, both MPPI and SVMPC fail for the static obstacles case.
@@ VIII-C Robot Manipulator on Surface
While obeying the table constraint the robot must also avoid 3 obstacles from the Yale-CMU-Berkeley (YCB) object dataset [65]. We enforce this with a constraint that the signed distance to the obstacles must be positive, which we compute from the meshes of the objects. Since signed distance functions (SDFs) are composable via the min operator, we combine the SDFs of the three obstacles into a single inequality constraint per timestep. To evaluate this constraint, offline we generate points on the surface of the robot. Online, we use forward kinematics to map all of these points to the world frame and evaluate their SDF value, selecting the minimum SDF value as the value of the constraint. To compute the gradient of the constraint, consider that for any surface point we can compute the gradient of the SDF value with respect to the point from the object mesh. We then use automatic differentiation to backpropagate this gradient through the forward kinematics to compute a gradient of the SDF value with respect to the joint configuration. Finally, to calculate an overall gradient, we use a weighted combination of the gradients for each surface point, with the weight computed via a softmin operation on the SDF values.
@@ VIII-D Robot Manipulator using wrench
We run this experiment for 20 trials with random initializations and show the results in the bottom row of Figure 10. This problem is challenging because the dynamics are based on a simple inaccurate geometric model. Compliance in the gripper causes deviation from this geometric model, and the model is only accurate so long as all constraints hold. . Our results show that CSVTO can succeed in all 20 trials with a goal threshold of 0.06 radians and achieves the lowest constraint violation. The next closest baseline, SVMPC-grad with $\lambda=10000$ succeeds 19/20 times with a goal threshold of 0.09 radians, dropping to 11/20 at 0.06 radians. We find that running IPOPT to convergence leads to poor performance, as the solver is unable to converge to a feasible solution. Limiting the maximum iterations to 200 for the initial warm-start and 20 for subsequent online iterations leads to improved task performance, achieving a success rate of 12/20.
We also demonstrate CSVTO on real hardware for the robot manipulator manipulating a wrench task, shown in Figure 1. The configuration-space trajectories generated by CSVTO are tracked using an impedance controller. We use the same hyperparameters as those in the simulator for this experiment. During execution, we applied disturbances by perturbing the robot end-effector. The impedance controller can reject small disturbances, but larger disturbances require re-planning from the perturbed location. Figure 1 shows one such perturbation. Despite large disturbances, our method was able to readjust the grasp and complete the task successfully.
@@ VIII-E Computation Time
The average computation time of CSVTO compared to baselines for all quadrotor experiments is shown in Table II. For this experiment, computing the gradient was a major computational bottleneck, thus for the sample-based methods we allowed them more iterations. We see that MPPI and SVMPC are faster than CSVTO with online trajectory computation times of $0.366s$, $0.439s$ and $0.589s$, respectively. For the no-obstacles and dynamic-obstacle cases, IPOPT is also faster than CSVTO with an average online computation time of $0.429s$ and $0.479s$ due to early termination. However, for the static obstacles case, this rises to $0.768s$ compared to CSVTO at $0.650s$. When running IPOPT to convergence, the solving time is substantially larger, with an average computation time for the static obstacle case of $15.8s$. We also see that the standard deviations are very large, due to the variability in how quickly the solver converges. Combining these with the results from Section VIII-B, we see that CSVTO outperforms IPOPT to convergence with substantially faster computational times.
@@ VIII-E2 Robot 7DoF Manipulator
The computational times for all methods on both 7DoF manipulation experiments are shown in Table III. For the manipulator on a surface experiment, the difference in computation speed of the sample-based vs gradient-based algorithms per iteration was less pronounced than for the quadrotor experiment. We thus kept the number of iterations the same for all experiments, with 100 warm-up iterations and 10 online iterations. CSVTO and IPOPT have similar computation times at $1.12s$ and $1.14s$ to compute a trajectory online. MPPI is again the fastest algorithm at $0.691s$ to generate a trajectory online, though the performance is lower both in terms of task success and constraint violation. Initial attempts to run IPOPT with a maximum of 1000 iterations took several minutes to solve, which rendered it impractical.
@@
Figure 9: The robot manipulator turning a wrench experimental set-up. The goal is to turn the wrench by 90 degrees. End effector path visualized for three different initial trajectories generated by Top) CSVTO, end-effector path traces an arc around the wrench center to turn the wrench. Bottom) IPOPT, generated trajectories are often poor, exhibiting very large step-sizes and non-smooth trajectories.
Figure 10: Results for the robot manipulator using the wrench. Left column shows success rate as we increase the size of the goal region. Right column shows the average constraint violation as a function of time time, where we compute the constraint violation at a given time via the maximum violation among the equality constraints.
@@ IX-C Limitations & Future Work
Our approach does have several limitations which we discuss here. First, our method requires that all costs and constraints are differentiable. This is a restrictive assumption, particularly when treating dynamics as a constraint. Many contact-rich robot manipulation tasks exhibit discontinuities that invalidate this assumption.
Second, our approach converts inequality constraints to equality constraints by introducing slack variables. While this is a natural way of incorporating inequality constraints into our method, it results in increasing the number of decision variables by the number of inequality constraints. This is likely to be problematic for long-horizon planning tasks with many inequality constraints. A possible solution would be solving a QP subproblem at every iteration to determine the active inequality constraints as in [49], however, this has the issue that we would need to solve an individual QP subproblem for every particle.
Finally, while our approach decomposes the kernel into a sum of kernels operating on sub-trajectories, each of these kernels is an RBF kernel. While the RBF has attractive properties, such as strict positive-definiteness and smoothness, we believe that exploring task-specific kernels for trajectory optimization is an interesting avenue for future work.
@@ X Conclusion
In this article, we presented Constrained Stein Variational Trajectory Optimization (CSVTO), an algorithm for performing constrained trajectory optimization on a set of trajectories in parallel. To develop CSVTO we formulated constrained trajectory optimization as a Bayesian inference problem, and proposed a constrained Stein Variational Gradient Descent (SVGD) algorithm inspired by O-SVGD [13] for approximating the posterior over trajectories. Our results demonstrate that CSVTO outperforms baselines in challenging highly-constrained tasks, such as a 7DoF wrench manipulation task, where CSVTO succeeds in 20/20 trials vs 12/20 for IPOPT. Additionally, our results demonstrate that generating diverse constraint-satisfying trajectories improves robustness to disturbances, such as moving obstacles, as well as robustness to initialization. [Matrix Derivative of $P(\tau)$] In Equation (32) we showed that the repulsive gradient is split into two terms, one of which contains the matrix derivative $\nabla_{[\tau]_{k}}P(\tau)$. In this section we show how to compute this derivative. For notational convenience, let $\tau\in\mathbb{R}^{N}$ (thus $P(\tau)\in\mathbb{R}^{NxN}$), $h(\tau)\in\mathbb{R}^{M}$ (where $M$ is the number of constraints), and we omit the dependence on $\tau$ when writing the constraint derivative $\nabla h(\tau)$. $\nabla_{[\tau]_{k}}P(\tau)$ is a matrix of shape $N\times N$. We refer to the second derivative of the $l$th constraint $\nabla^{2}h_{l}(\tau)$ as $H_{l}$, which is an $N\times N$ matrix. The matrix derivative $\nabla_{[\tau]_{k}}P(\tau)$, as defined in Equation 32, can be expanded into three terms:

### v3

@@ VII-A7 Initialization
As introduced in section IV, we have a user-specified prior on controls $p(U)$. To initialize CSVTO on a new problem, we proceed by sampling from this prior $p(U)$ and using the dynamics $f(x_{t},u_{t})$ to generate sampled trajectories. In this way, we ensure that the initial trajectory satisfies the dynamics constraints.
We use a different initialization scheme when running trajectory optimization online in a receding horizon fashion as in Algorithm 2, as it is typical to warm-start the optimization with the solution from the previous timestep. For a single particle, the trajectory consists of $\tau=(\mathbf{x}_{1},...,\mathbf{x}_{T},\mathbf{u}_{0},...,\mathbf{u}_{T-1)})$. The shift operation computes $\tau^{\prime}=(\mathbf{x}_{2},...,\mathbf{x}^{\prime}_{T+1},\mathbf{u}_{1},...,\mathbf{u}^{\prime}_{T})$. Here $\mathbf{x}^{\prime}_{T+1},\mathbf{u}^{\prime}_{T}$ is the initialization for the newly considered future timestep. The initializations $\mathbf{x}^{\prime}_{T+1},\mathbf{u}^{\prime}_{T}$ may be chosen in a problem-specific way. In our approach, we choose them by duplicating the previous timestep’s state and control, i.e. $(\mathbf{x}^{\prime}_{T+1},\mathbf{u}^{\prime}_{T})=(\mathbf{x}_{T},\mathbf{u}_{T-1})$.
When running the algorithm with inequality constraints, for both the online and warm-start optimizations we initialize the slack variable $\mathbf{z}$ with $\mathbf{z}=\sqrt{2|g(\tau)|}$ so that trajectories satisfying the inequality constraint are initialized to satisfy the transformed equality constraint.
@@ VII-A8 Re-sampling
$w_{i}=\frac{\exp(-\frac{\hat{C}_{\lambda}(\hat{\tau}_{i})}{\beta})}{\sum^{N}_{j}\exp(-\frac{\hat{C}_{\lambda}(\hat{\tau}_{j})}{\beta})},$ (51)
@@
1:function CSVTO($\mathbf{x}_{0},\tau,K,\texttt{anneal}$)
2:     $\mathbf{z}\leftarrow\sqrt{2|g(\tau)|}$
3:     $\hat{\tau}\leftarrow[\tau,\mathbf{z}]^{T}$
@@
14:         $\mathbf{u}_{t}\leftarrow\tau*$
15:         $\mathbf{x}_{t}\leftarrow$ StepEnv$(\mathbf{u}_{t})$
@@ VIII-A Baselines
We evaluate our method on a 12DoF underactuated quadrotor problem. The goal is to navigate the quadrotor from a start state to a goal state. We chose this problem to demonstrate our approach on a problem with complex nonlinear underactuated dynamics. The experimental setup is shown in Figures 3 and 4. The state of the quadrotor is $\mathbf{x}=[x,y,z,p,q,r,\dot{x},\dot{y},\dot{z},\dot{p},\dot{q},\dot{r}]^{T}$, where $(x,y,z)$ is the 3D position and $(p,q,r)$ are the Euler angles. The control is the thrust $\mathbf{u}=[u_{1},u_{2},u_{3},u_{4}]^{T}\in\mathbb{R}^{4}$. We place bounds constraints on the $(x,y)$ location of the quadrotor to be within a $10m\times 10m$ area centred at $(0,0)$. The goal is to travel from start locations sampled uniformly from $x,y\in[3.0m,4.5m]$ to a goal location of $(4,4)$ within 100 time steps. We place an equality constraint that the quadrotor must travel along a nonlinear surface $z=f_{surf}(x,y)$. For this surface, we sample $z$ values from a Gaussian Process (GP) prior with an RBF kernel and zero mean function on a $10\times 10$ grid of $(x,y)$ points. We use the sampled values as observations for a GP with the same kernel and mean function and use the corresponding posterior mean function as the equality constraint. We sample a single surface in this way and use it for all experiments. The dynamics for the 12DoF quadrotor are from [71] and are given by
@@
where $c(p),s(p),t(p)$ are $\cos,\sin,\tan$ functions, respectively. We use parameters $m=1kg,I_{x}=0.5kg\cdot m^{2},I_{y}=0.1kg\cdot m^{2},I_{z}=0.3kg\cdot m^{2},K=5,g=-9.81m\cdot s^{-2}$. We use the same dynamics both for planning and for simulation.
We consider three variants of this task with different obstacle avoidance constraints: 1) We consider the case with no obstacles; 2) We consider the case of static obstacles. For the static obstacles case, we wish to demonstrate our method in a cluttered environment with arbitrarily shaped obstacles. We do this by generating the obstacles similarly to the surface constraint, which results in a smooth obstacle constraint. We consider a constraint function $f_{obs}(x,y)$, where the obstacle-free region is $\{(x,y),f_{obs}(x,y)\leq 0\}$. We sample values for $f_{obs}(x,y)$ from a GP prior with an RBF kernel and a constant mean function of $-0.5$, so that there is a bias towards being obstacle-free, on a $10\times 10$ grid of $(x,y)$ points. We then use these points as the observations for a GP with the same mean function and kernel as the GP prior. We also add observations at $(-4,-4)$ and $(4,4)$ of $-2$, to ensure the start and goal regions are obstacle free. We use the resulting GP posterior mean function as $f_{obs}(x,y)$. We do this once and keep the same obstacle constraint for all trials. The resulting obstacle constraint is shown in Figure 4; 3) Finally, we consider a cylindrical obstacle in the x-y plane that moves during the trial in a path that is unknown to the planner; at every timestep, the planner plans assuming the obstacle will remain fixed. If the quadrotor collides with an obstacle during execution then we consider the task failed.
@@
The control cost is equivalent to the prior on controls $p(u_{t})=\mathcal{N}(0,2R^{-1})$. The values we use for the costs are
$Q$ $=\text{Diag}(5,5,0.5,2.5,2.5,0.025,1.25,1.25,1.25,2.5,2.5,2.5)$ (55)
@@
We ran the quadrotor experiments for the three different obstacle cases for 20 trials with randomly sampled starts. The results are shown in Figure 6. CSVTO succeeds for 20/20 trials for the no obstacles and dynamic obstacles cases, and 19/20 for the static obstacle case, all with a goal threshold of 0.3m. For the static-obstacle and dynamic-obstacle experiments, IPOPT is the next best performing with 20/20 trials for no-obstacles at a goal threshold of 0.4m, but success falls to 15/20 for both the static-obstacles and dynamic-obstacle case. We see that running IPOPT with more iterations improves performance for the static obstacles case, but in the other two cases, there is no significant difference in performance when allowing IPOPT to run until convergence. However, running IPOPT to convergence has substantially higher computation time, which we will discuss further in section VIII-E. For the no-obstacles and dynamic-obstacle cases, we see that sample-based methods perform well according to the task success rate, however, they fail to satisfy the surface equality constraint. In addition, both MPPI and SVMPC fail for the static obstacles case.
@@ VIII-C Robot Manipulator on Surface
While obeying the table constraint the robot must also avoid 3 obstacles from the Yale-CMU-Berkeley (YCB) object dataset [72]. We enforce this with a constraint that the signed distance to the obstacles must be positive, which we compute from the meshes of the objects. Since signed distance functions (SDFs) are composable via the min operator, we combine the SDFs of the three obstacles into a single inequality constraint per timestep rather than an inequality constraint per obstacle. This is to reduce the total number of inequality constraints, as introducing more inequality constraints results in more slack variables and a higher dimensional problem. To evaluate this constraint, offline we generate points on the surface of the robot. Online, we use forward kinematics to map all of these points to the world frame and evaluate their SDF value, selecting the minimum SDF value as the value of the constraint. To compute the gradient of the constraint, consider that for any surface point we can compute the gradient of the SDF value with respect to the point from the object mesh. We then use automatic differentiation to backpropagate this gradient through the forward kinematics to compute a gradient of the SDF value with respect to the joint configuration. Finally, to calculate an overall gradient, we use a weighted combination of the gradients for each surface point, with the weight computed via a softmin operation on the SDF values.
@@ VIII-D Robot Manipulator using wrench
We run this experiment for 20 trials with random initializations and show the results in the bottom row of Figure 10. This problem is challenging because the dynamics are based on a simple inaccurate geometric model. Compliance in the gripper causes deviation from this geometric model, and the model is only accurate so long as all constraints hold. Our results show that CSVTO can succeed in all 20 trials with a goal threshold of 0.06 radians and achieves the lowest constraint violation. The next closest baseline, SVMPC-grad with $\lambda=10000$ succeeds 19/20 times with a goal threshold of 0.09 radians, dropping to 11/20 at 0.06 radians. We find that running IPOPT to convergence leads to poor performance, as the solver is unable to converge to a feasible solution. Limiting the maximum iterations to 200 for the initial warm-start and 20 for subsequent online iterations leads to improved task performance, achieving a success rate of 12/20.
We also demonstrate CSVTO on real hardware for the robot manipulator manipulating a wrench task, shown in Figure 1. After generating a configuration-space trajectory using CSVTO, we command the robot to move to the first configuration waypoint of that trajectory using a joint impedance controller. Once the robot has reached the desired waypoint, we perform re-planning to generate a new configuration-space trajectory. We use the same hyperparameters as those in the simulator for this experiment. During execution, we applied disturbances by perturbing the robot end-effector. The impedance controller can reject small disturbances, but larger disturbances require re-planning from the perturbed location. Figure 1 shows one such perturbation. Despite large disturbances, our method was able to readjust the grasp and complete the task successfully.
@@ VIII-E Computation Time
The average computation time of CSVTO compared to baselines for all quadrotor experiments is shown in Table II. For this experiment, computing the gradient was a major computational bottleneck, thus for the sample-based methods we allowed them more iterations. We see that MPPI and SVMPC are faster than CSVTO with online trajectory computation times of $0.366s$, $0.439s$, and $0.589s$, respectively. For the no-obstacles and dynamic-obstacle cases, IPOPT is also faster than CSVTO with an average online computation time of $0.429s$ and $0.479s$ due to early termination. However, for the static obstacles case, this rises to $0.768s$ compared to CSVTO at $0.650s$. When running IPOPT to convergence, the solving time is substantially larger, with an average computation time for the static obstacle case of $15.8s$. We also see that the standard deviations are very large, due to the variability in how quickly the solver converges. Combining these with the results from Section VIII-B, we see that CSVTO outperforms IPOPT to convergence with substantially faster computation times.
@@ VIII-E2 Robot 7DoF Manipulator
The computation times for all methods on both 7DoF manipulation experiments are shown in Table III. For the manipulator on a surface experiment, the difference in computation speed of the sample-based vs gradient-based algorithms per iteration was less pronounced than for the quadrotor experiment. We thus kept the number of iterations the same for all experiments, with 100 warm-up iterations and 10 online iterations. CSVTO and IPOPT have similar computation times at $1.12s$ and $1.14s$ to compute a trajectory online. MPPI is again the fastest algorithm at $0.691s$ to generate a trajectory online, though the performance is lower both in terms of task success and constraint violation. Initial attempts to run IPOPT with a maximum of 1000 iterations took several minutes to solve, which rendered it impractical.
@@
Figure 9: The robot manipulator turning a wrench experimental set-up. The goal is to turn the wrench by 90 degrees. End-effector planned path at the first time-step visualized for three different initial trajectories generated by CSVTO (Top) and IPOPT (Bottom). CSVTO’s end-effector path traces an arc around the wrench center to turn the wrench, while IPOPT paths are often poor, containing very large steps and lacking smoothness
Figure 10: Results for the robot manipulator using the wrench. Left column shows the success rate as we increase the size of the goal region. The right column shows the average constraint violation as a function of time time, where we compute the constraint violation at a given time via the maximum violation among the equality constraints.
@@ IX-C Limitations & Future Work
Differentiability
Our method requires that all costs and constraints are differentiable. This is a restrictive assumption, particularly when treating dynamics as a constraint. Many contact-rich robot manipulation tasks exhibit discontinuities that invalidate this assumption.
Slack variables
Our approach converts inequality constraints to equality constraints by introducing slack variables. While this is a natural way of incorporating inequality constraints into our method, it results in increasing the number of decision variables by the number of inequality constraints. This is likely to be problematic for long-horizon planning tasks with many inequality constraints. A possible solution would be solving a QP subproblem at every iteration to determine the active inequality constraints as in [56], however, this has the issue that we would need to solve an individual QP subproblem for every particle.
Computation time inadequate for real-time control
We note from Table II, in the dynamic obstacle quadrotor task the average computation time for online trajectory generation is $0.643s$ for CSVTO, compared to MPPI, the fastest baseline, taking $0.373s$. In this case, the solve times for the current implementation of CSVTO and all baselines are insufficient for real-time reactive control. Our method, all baselines other than IPOPT, and all cost and constraint functions were implemented in Python, using automatic differentiation in PyTorch to compute the relevant first and second derivatives. Implementing these methods in C++, using a library such as CasADI [73] for automatic differentiation, may enable real-time performance on these systems in future work.
Kernel selection
While our approach decomposes the kernel into a sum of kernels operating on sub-trajectories, each of these kernels is an RBF kernel. While the RBF has attractive properties, such as strict positive-definiteness and smoothness, we believe that exploring task-specific kernels for trajectory optimization is an interesting avenue for future work.
@@ X Conclusion
In this article, we presented Constrained Stein Variational Trajectory Optimization (CSVTO), an algorithm for performing constrained trajectory optimization on a set of trajectories in parallel. To develop CSVTO we formulated constrained trajectory optimization as a Bayesian inference problem, and proposed a constrained Stein Variational Gradient Descent (SVGD) algorithm inspired by O-SVGD [23] for approximating the posterior over trajectories with a set of particles. Our results demonstrate that CSVTO outperforms baselines in challenging highly-constrained tasks, such as a 7DoF wrench manipulation task, where CSVTO succeeds in 20/20 trials vs 12/20 for IPOPT. Additionally, our results demonstrate that generating diverse constraint-satisfying trajectories improves robustness to disturbances, such as changes in the environment, as well as robustness to initialization.
[Matrix Derivative of $P(\tau)$] In Equation (33) we showed that the repulsive gradient is split into two terms, one of which contains the matrix derivative $\nabla_{[\tau]_{k}}P(\tau)$. In this section, we show how to compute this derivative. For notational convenience, let $\tau\in\mathbb{R}^{N}$ (thus $P(\tau)\in\mathbb{R}^{N\times N}$), $h(\tau)\in\mathbb{R}^{M}$ (where $M$ is the number of constraints), and we omit the dependence on $\tau$ when writing the constraint derivative $\nabla h(\tau)$. $\nabla_{[\tau]_{k}}P(\tau)$ is a matrix of shape $N\times N$. We refer to the second derivative of the $l$th constraint $\nabla^{2}h_{l}(\tau)$ as $H_{l}$, which is an $N\times N$ matrix. The matrix derivative $\nabla_{[\tau]_{k}}P(\tau)$, as defined in Equation (33), can be expanded into three terms: